Daily demand for gasoline at Billy-Bob's Mobile Station is described by D(p) = 980-300p, where p is the price in dollars. Billy-Bob's supply is S(p) =-2980 + 3000p.Suppose that the state government places a tax of 18 cents on every gallon of gasoline sold. What is the deadweight loss resulting from the tax?
Answer #1

(), - а - Зеер — When peo, O = 980. When Oro, P = $3.26 ( 4 -21 + Зээр Air C, D, P. p. 49 Equilibrium ocurs when ad og 9 съ -Os - 2980 + 3000 (P - 0.18) = -298 + 3000 P - 5% Jos = -3520 + 3000 p Now, equate of and Os, we getre 980-3 oo P = -352 +3000Price supply after tax 3.26 Supply test toz to18 Demon 0.99 572 620 980 980 Quantity dax Deadweight loss resulting from the =